OpenStudy (anonymous):

please hellp i'm begging Find all of the zeros of the equations 4) x3 – 5x2 + 6x = 0 5) x4 – 8x2 – 9 = 0 6) 5x3 – 5x = 0

OpenStudy (mathmale):

(4) has x in each term of the left side, so factor out x. What's left? How would you go about factoring x^2 - 5x + 6?

OpenStudy (anonymous):

what is your question exactly

OpenStudy (mathmale):

I'm asking you to factor x out of x^ – 5x^2 + 6x = 0. You'll be left with a factor that's a quadratic expression. How would you factor that expression?

OpenStudy (anonymous):

um i don't know :'( i take cyber school and math is the hardest thing ever online

OpenStudy (whpalmer4):

\[x^3-5x^2+6x = 0\]All of the terms have \(x\) in them, so you can factor out \(x\). What do you get when you do that?

OpenStudy (anonymous):

(x-3)(x-2)? i have no idea

OpenStudy (whpalmer4):

What is \(x^3/x\)? What is \(-5x^2/x\)? What is \(6x/x\)?

OpenStudy (anonymous):

whpalmer will help you pkay

OpenStudy (anonymous):

wait how do i figure that out?

OpenStudy (anonymous):

ask him.. i have no idea etheir

OpenStudy (mathmale):

\[\frac{ x^3−5x^2+6x=0 }{ x }\] may be a little easier to understand.

OpenStudy (whpalmer4):

If you have \(xa + xb + xc\), that's the same as \(x(a+b+c)\), right?

OpenStudy (mathmale):

Here you are asked to divide each term of the equation by x. \[\frac{ x ^{3} }{ x }=x ^2\]. This is precisely what whpalmer4 is asking you to do.

OpenStudy (anonymous):

exponents confuse me

OpenStudy (whpalmer4):

You can write that polynomial as\[x*x^2 - x*5x+x*6\]or\[x(x^2-5x+6)\] exponents are just a shorthand. \[x^2 = x*x\]\[x^3 = x*x*x\]\[x^1 = x\]The only moderately tricky one is \[x^0 = 1, \text{ as long as }x\ne0\]

OpenStudy (whpalmer4):

So the division of \[\frac{x^3}{x} = x^2\] is the same as \[\frac{x*x*x}{x} = \frac{\cancel{x*}x*x}{\cancel{x}} = x*x = x^2\]

OpenStudy (anonymous):

ohh okay

OpenStudy (whpalmer4):

So, we can factor out that \(x\) from the polynomial: \[x^3-5x^2 + 6x = x(x^2-5x+6)\] Now we want to factor \(x^2-5x+6\) We'll have two product terms: \[(x+a)(x+b) = x*x + b*x + a*x + a*b = x^2 + (a+b)x + a*b\] Now we need to pick two numbers \(a,b\) to make \[x^2-5x+6\]match with \[x^2 + (a+b)x + a*b\] For that to be true, we need \(a+b = -5\) and \(a*b = 6\) Can you think of a pair of numbers that makes that true?

OpenStudy (anonymous):

that negative throws me off because it could be a= 1 b=-6 but then when you multiply it wouldn't it make 6 negative?

OpenStudy (whpalmer4):

here are the ways we can get a*b = 6 1*6 2*3 3*2 6*1 -1*-6 -2*-3 -3*-2 -6*-1 do any of those pairs add up to -5?

OpenStudy (anonymous):

the -2 and -3 ?

OpenStudy (anonymous):

i'm sorry it is midnight here and my brain is not working haha

OpenStudy (whpalmer4):

right, you got it! \[(x-2)(x-3) = x*x - 2x - 3x -2(-3) = x^2 - 5x + 6\] So the first equation factors to \[x(x-2)(x-3) = 0\] What does that mean? What are the values of \(x\) that can make that true?

OpenStudy (anonymous):

hmm

OpenStudy (whpalmer4):

Well, if I tell you that I've multiplied two numbers together, and the result is 0, does that suggest anything about either number?

OpenStudy (anonymous):

one of them is 0

OpenStudy (whpalmer4):

Bingo!!! So, if we have three numbers multiplied together, and the result is 0, that means that any one of the 3 might be 0, right?

OpenStudy (mathmale):

WHP: Great work!!

OpenStudy (whpalmer4):

\[x(x-2)(x-3)= 0\]means that either \[x=0\]\[x-2=0\]\[x-3=0\]What are the values of \(x\) (the zeros) which make those equations true? The first one is pretty easy :-)

OpenStudy (anonymous):

wait so i'm just plugging in numbers to make them equal zero?

OpenStudy (whpalmer4):

or solving. \[x = 0\]Well, that's easy to solve :-) \[x - 2 = 0\]add 2 to both sides\[x-2+2 = 0+2\]\[x = 2\] \[x-3=0\]add 3 to both sides\[x-3+3 = 0+3\]\[x=3\] So x = 0, x = 2, x = 3 are all zeros of that function, because they make the value of the function be 0. For example: \[x^3-5x^2+6x = 0\]Plug in x = 0:\[(0)^3-5(0)^2+6(0) = 0-0 + 0 = 0\]Plug in x = 2:\[(2)^3-5(2)^2 + 6(2) = 2*2*2-5*2*2+6*2 \]\[= 8 - 20 + 12 = 8+12 - 20 = 20 - 20 = 0\] Plug in x = 3:\[(3)^3 - 5(3)^2+6(3) = 3*3*3 - 5*3*3 + 6*3 = 27-45+18 = 0\] So those are the zeros of this function. Make sense?

OpenStudy (anonymous):

pretty much yeah

OpenStudy (anonymous):

you are like the smartest person ever haha

OpenStudy (whpalmer4):

I'm guessing you might need help with the other two... \[x^4 – 8x^2 – 9 = 0\] Here we need two numbers that multiply to -9, and add to -8. What could they be?

OpenStudy (whpalmer4):

I guess you called it a night. Sounds like a good plan.